Doc. Math. J. DMV 227

Section 4 Algebraic Geometry

In case of several authors, Invited Speakers are marked with a ^{∗}.

Paul S. Aspinwall: String Theory and Duality . . . . . . . . . . . . . . . . . . . . II 229
Victor V. Batyrev: Mirror Symmetry and Toric Geometry . . . . . . . . II 239
Maurizio Cornalba: Cohomology of Moduli Spaces of Stable Curves II 249
A. J. de Jong: Barsotti-Tate Groups and Crystals . . . . . . . . . . . . . . . . . . II 259
Mark L. Green: Higher Abel-Jacobi Maps . . . . . . . . . . . . . . . . . . . . . . . . . II 267
M. Kapranov: Operads and Algebraic Geometry . . . . . . . . . . . . . . . . . . . II 277

Documenta Mathematica · Extra Volume ICM 1998 · II ·

228

Documenta Mathematica · Extra Volume ICM 1998 · II ·